OpenStudy (anonymous):

Jane knits 72 stitches to the line and uses 1/4 inch of yarn in each of the stitches. How many lines can she knit with 10 yards of yarn? (1 yard = 3feet; 1 foot =12 inches) A. 10 B. 20 C. 30 D. 40 E. 50

OpenStudy (nurali):

given total 72 stitches and 1 stitch takes 1/4 inch, so total yarn used : 18 inch 36 inch = 3 feet = 1 yard divide by 2 18 inch = 3/2 feet = 1/2 yard 1/2 yard ---1 line 10 yard -- 20 lines